OpenStudy (anonymous):

Find how many quarts of 6% butterfat milk and 2% butterfat milk should be mixed to yield 60 quarts of 4% butterfat milk.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (xapproachesinfinity):

where is the rest of the problem?

OpenStudy (anonymous):

The mixture should contain _ quarts of the 6 percent butterfat milk

OpenStudy (aum):

Assume "x" quarts of 6% butterfat milk and (60-x) quarts of 2% butterfat milk are mixed to yield 60 quarts of 4% butterfat milk. The fat content in "x" quarts of 6% butterfat milk = x * 6/100 = 0.06x The fat content in (60-x) quarts of 2% butterfat milk = (60-x) * 2/100 = 0.02(60-x) Total fat content = 0.06x + 0.02(60-x) = 0.06x + 1.2 - 0.02x = 0.04x + 1.2 The fat content in 60 quarts of 4% butterfat milk = 60 * 4/100 = 2.4 Therefore, 0.04x + 1.2 = 2.4 solve for x.